OpenStudy (anonymous):

What is the vertex form of the equation? Y=^2+4x-3 A. Y=(x-2)^2-7 B. Y=(x+2)^2-2 C. Y=(x-2)^2+7 D. Y=(x+2)^2+7

OpenStudy (anonymous):

*x^2

ganeshie8 (ganeshie8):

\(\large x^2+4x -3\)

ganeshie8 (ganeshie8):

do u knw the vertex form ?

OpenStudy (anonymous):

No I'm confused

ganeshie8 (ganeshie8):

vertex form :- \(\large a(x-h)^2 + k\)

OpenStudy (anonymous):

Okay I get that but how do you get the answer

ganeshie8 (ganeshie8):

\(\large x^2+4x -3 \) \(\large x^2+2 \times 2 x -3 \)

ganeshie8 (ganeshie8):

\(\large x^2+4x -3 \) \(\large x^2+2 \times 2 x -3 \) \(\large x^2+2 \times 2 x + 4 - 4 -3 \)

ganeshie8 (ganeshie8):

\(\large x^2+4x -3 \) \(\large x^2+2 \times 2 x -3 \) \(\large x^2+2 \times 2 x + 2^2 - 4 -3 \)

ganeshie8 (ganeshie8):

\(\large x^2+4x -3 \) \(\large x^2+2 \times 2 x -3 \) \(\large x^2+2 \times 2 x + 2^2 - 4 -3 \) \(\large (x+2)^2 - 7 \)

OpenStudy (anonymous):

I okay thank you now I get it :)

ganeshie8 (ganeshie8):

cool :) but looks thats not there in ur options ! options might be incorrect, or the given equation...

OpenStudy (anonymous):

I think it's b because I by accidentally out 2

ganeshie8 (ganeshie8):

ohk... :)